OpenStudy (anonymous):

Which of the following are solutions to the equation below?(2x + 3)2 = 1

OpenStudy (shubhamsrg):

either take sqrt of both sides or expand LHS and form quadratic eqn,,both will yield you same 2 answers..

OpenStudy (shubhamsrg):

altough with a bit of thinking,,you should be able to do the ques orally,,just saying ;)

OpenStudy (anonymous):

LOL....alright ill attempt

jhonyy9 (jhonyy9):

2x+3=+/- 1 2x+3=1 2x=-2 x=-2/2 =-1 and 2x+3=-1 2x=-4 x=-4/2 =-2 x_1=-1 x_2=-2